graph the equation and idenify the y intercept 6y+5x=-18
Answer by wuwei96815(245)

6y+5x=-18
To find the y-intercept let x=0
6y+5(0)=-18
6y=-18
y=-3
The y-intercept is (0,-3)

To find the x-intrcept let y=0
6(0)+5x=-18
5x=-18
x=-18/5
The x-intrecept is (-18/5,0)

The line 6y+5x=-18 will pass through the points (0,-3) and (-18/5,0)